# Compression of Higher Order Ambisonics with Multichannel RVQGAN

### v2

@@ Abstract
A multichannel extension to the RVQGAN neural coding method is proposed, and realized for data-driven compression of third-order Ambisonics audio. The input- and output layers of the generator and discriminator models are modified to accept multiple (16) channels without increasing the model bitrate. We also propose a loss function for accounting for spatial perception in immersive reproduction, and transfer learning from single-channel models. Listening test results with 7.1.4 immersive playback show that the proposed extension is suitable for coding scene-based, 16-channel Ambisonics content with good quality at 16 kbps.
@@ 2 Background
The majority of neural audio coding research has been focused on single-channel aka mono audio compression. However, most modern audio content has more than one channel, making relevant the spatial and immersive aspects of audio. In recent decades, formats and reproduction systems utilizing more than two channels have emerged, starting from 5.1 surround and culminating with the immersive audio formats of recent years [8, 9, 10, 11, 12]. To our knowledge, the few attempts to apply neural coding methods to multichannel audio coding amount to processing a stereo input as dual-mono, thereby doubling the bitrate [13]. These works do not take advantage of the interchannel correlations present in the data.
@@ 4 Results and Discussion
Although not explicitly tested here, the present modeling principle is applicable to multichannel audio formats other than 3rd-order Ambisonics (e.g. 5.1, 7.1, 7.1.4 etc.), as long as there is sufficient training data. Despite the promising results, the main caveats of the present method, and most neural codecs remain 1) the lack of scaling to different bitrates without retraining or model modification, and 2) the model computational complexity. For these, future work will hopefully provide solutions.

### v3

@@ Abstract
A multichannel extension to the RVQGAN neural coding method is proposed, and realized for data-driven compression of third-order Ambisonics audio. The input- and output layers of the generator and discriminator models are modified to accept multiple (16) channels without increasing the model bitrate. We also propose a loss function for accounting for spatial perception in immersive reproduction, and transfer learning from single-channel models. Listening test results with 7.1.4 immersive playback show that the proposed extension is suitable for coding scene-based, 16-channel Ambisonics content with good quality at 16 kbps when trained and tested on the EigenScape database. The model has potential applications for learning other types of content and multichannel formats.
@@ 2 Background
The majority of neural audio coding research has been focused on single-channel aka mono audio compression. However, most modern audio content has more than one channel, making relevant the spatial and immersive aspects of audio. In recent decades, formats and reproduction systems utilizing more than two channels have emerged, starting from 5.1 surround and culminating with the immersive audio formats of recent years [8, 9, 10, 11, 12]. Some attempts to apply neural coding methods to multichannel audio coding amount to processing a stereo input as dual-mono, thereby doubling the bitrate [13]. These works do not take advantage of the interchannel correlations present in the data. As an exception, Kleijn et al. [14] generated multichannel (in the experiment only stereo) as a mono channel + spatial metadata, and used generative methods for the spatial information. However, as opposed to being more generative in nature, the present study aims to replicate and compare to the original multichannel spatial impression.
@@ 4 Results and Discussion
As the main caveat, it should be noted that informal listening with few musical samples outside the EigenScape database indicated that applying the proposed model trained only on scene-based material may not be as suitable for all content. Other content types or more generic models can also require increased bitrate to achieve good quality. More general training and application of the model with e.g. musical and cinematic content is left for future.
Although not tested here, the present modeling principle is also applicable to multichannel audio formats other than 3rd-order Ambisonics (e.g. 5.1, 7.1, 7.1.4 etc.), as long as there is sufficient training data. Despite the promising results, the main caveats of the present method, and most neural codecs remain 1) the lack of scaling to different bitrates without retraining or model modification, and 2) the model computational complexity. For these, future work will hopefully provide solutions.